Use the distributiive law to help solve.
6(w-2)= -12
I do not understnad the distributive law what is it.

Answer by checkley77(12844)

6(W-2)=-12
6W-12=-12
6W=-12+12
6W=0
W=0 ANSWER.
PROOF
6(0-2)=-12
6*-2=-12
-12=-12